OpenStudy (anonymous):

Dachelle does pushups every morning. She does 3 sets of 8 pushups, 7 days a week. How many pushups does Dachelle do in one week?

OpenStudy (anonymous):

Number of pushups in a day= 3*8 = 24 Number of pushups in a week = 24*7 = 168